Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland Mark Dennehy,

DURC Secretary

Vision Lab

O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Simon Kennedy  14th March 2002

A Chara,

Simon Kennedy has been a long-standing member of the Dublin University Rifle Club in the past and now holds honorary member status. He has represented Ireland in competition outside the republic and is of good character.

Misé le meas,

Mark Dennehy, DURC Secretary